Evaluate:
a. 5! =
b. small (7) P small (5) 7P5 =
c. small (7) C cmall (5) 7C5 =

Answer by stanbon(75887)

Evaluate:
a. 5! = 5*4*3*2*1 = 120
------------------------------------
b. small (7) P small (5); 7P5 = 7!/[(7-5)!] = (7*6*5*4*3*2*1)/(2*1)
= 7*6*5*4*3 = 2520
----------------------------------
c. small (7) C cmall (5); 7C5 = (7*6*5*4*3)/(1*2*3*4*5) = (7*6)/(1*2)= 21
=================================
Cheers,
Stan H.